In 10 years, a child’s age will be 6 years less than triple her current age. Find the child’s current age

3X-6X + 10 -lg $6-10 46 The childs comment
Answer #1

Let the child current age = le After ten year his age will be = 10+ 0 According to question 10 + x = 3x - 6 - 3*-* = 10+6 280